OpenStudy (anonymous):

find f when gradient f =2xyi+x^2j

OpenStudy (anonymous):

@JamesJ can u plz come help me?

OpenStudy (anonymous):

James ur pic looks cool with a purple border i love it

OpenStudy (jamesj):

First, what's the definition of gradient? In other words, suppose F is a scalar field, F = F(x,y), then what is grad F?

OpenStudy (anonymous):

grad F = dF/dxi+dF/dyj

OpenStudy (jamesj):

Right, where d here is partial derivative. So that means you know that dF/dx = 2xy dF/dy = x^2 Now, given that, what function F satisfies those equations?

OpenStudy (anonymous):

f=yx^2

OpenStudy (anonymous):

uughhh now i see it

OpenStudy (jamesj):

f(x,y) = x^2y + C yes

OpenStudy (anonymous):

idk y i cldnt see it b4

OpenStudy (jamesj):

np

OpenStudy (anonymous):

Thanks :)